Barney is flying a 1200kg helicopter. If he is accelerating towards the ground at 3.15m/s^2, what is the force applied to the rotors by the air?
1
Expert's answer
2020-04-05T14:43:48-0400

As per the given question,

Mass of the helicopter =1200 kg

downward acceleration "(a)=3.15 m\/sec^2"

gravitational acceleration "=9.8 m\/sec^2"

Let the upward force is F.

So, "mg-F= ma"

"\Rightarrow F=mg-ma"

"\Rightarrow F=1200(9.8-3.15)=6.65\times 1200N"

"\Rightarrow F=7980N"